OpenStudy (anonymous):

please someone help ill give fan and medal Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@Preetha @pooja195 @Hero please help

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

Can you help @Loser66 ?